An airplane takes 9 hours to travel a distance of 6480km against the wind. The return trip takes 8 hours with the wind. What is the rate of the plane in still air and what is the rate of the wind?
Answer by checkley77(12844)

D=RT
6,480=R9
R=6,480/9=720 KMH AGAINST THE WIND.
6,480=R8
R=6,480/8
R=810 KMH WITH THE WIND.
(810-720)/2=90/2=45 KMH IS THE SPEED OF THE WIND.
PROOF:
720+45=765 SPEED OF THE PLANE IN STILL AIR
810-45=765 DITTO.